# DMDSpeech: Distilled Diffusion Model Surpassing The Teacher in Zero-shot Speech Synthesis via Direct Metric Optimization

### v1

@@ 4.4 Ablation Study
Effects of Distribution Matching Distillation Using a single sampling step resulted in significantly degraded performance compared to the full DMDSpeech model. While increasing to four sampling steps improved naturalness and sound quality to approach the teacher model’s level, speaker similarity remained significantly lower. Interestingly, the speaker verification model’s SIM score showed only a slight decrease, suggesting a phenomenon we term mode shrinkage (Figure 2), where distillation emphasizes high-probability regions of the data distribution. This focus can result in a more generic speaker profile, reducing perceived uniqueness in the prompt speaker’s voice, while maintaining global speaker features as reflected in the SIM score. To address this, we introduced speaker verification loss in this work to better capture the distinct characteristics of the prompt speaker.
@@
Table 4: Ablation study comparing our proposed model with different conditions. MOS-N, MOS-Q, SMOS-V, and SMOS-S are reported as mean ($\pm$ standard error). Models with statistically significant differences ($p<0.05$) compared to DMDSpeech are marked with one asterisk (*). Additional evaluation results on emotion reflection are presented in Table 7.
Model MOS-N MOS-Q SMOS-V SMOS-S WER SIM $\text{CV}_{f_{0}}$ Teacher (N=128) 4.22 ($\pm$ 0.04) 4.40 ($\pm$ 0.03) 4.03 ($\pm$ 0.04) 3.87 ($\pm$ 0.04) 9.51 0.55 0.70 DMD 2 only (N=1) 3.11 ($\pm$ 0.05) 2.99 ($\pm$ 0.05) 2.57 ($\pm$ 0.05) 2.74 ($\pm$ 0.05) 5.93 0.42 0.68 DMD 2 only (N=4) 4.19 ($\pm$ 0.03) 4.43 ($\pm$ 0.04) 3.69 ($\pm$ 0.05) 3.62 ($\pm$ 0.05) 5.67 0.53 0.61 $+\mathcal{L}_{\text{CTC}}$ only 4.25 ($\pm$ 0.04) 4.42 ($\pm$ 0.03) 3.73 ($\pm$ 0.05) 3.62 ($\pm$ 0.05) 1.79 0.55 0.57 $+\mathcal{L}_{\text{SV}}$ only 4.07 ($\pm$ 0.04) 4.33 ($\pm$ 0.03) 4.35 ($\pm$ 0.04) 4.15 ($\pm$ 0.04) 6.62 0.70 0.61 DMDSpeech (N=4) 4.27 ($\pm$ 0.03) 4.45 ($\pm$ 0.03) 4.35 ($\pm$ 0.03) 4.16 ($\pm$ 0.03) 1.94 0.69 0.58 B. S. $96\rightarrow 16$ 4.20 ($\pm$ 0.04) 4.30 ($\pm$ 0.03) 4.27 ($\pm$ 0.04) 4.11 ($\pm$ 0.04) 3.38 0.67 0.60
@@ 5 Conclusions
In this work, we presented DMDSpeech, a distilled diffusion-based text-to-speech (TTS) model based on prompt continuation. By employing distribution matching distillation (DMD), our model generates high-quality speech in just 4 steps and enables direct metric optimization. Through speaker verification (SV) and connectionist temporal classification (CTC) losses, DMDSpeech significantly improves speaker similarity and text-speech alignment, outperforming several state-of-the-art baselines.
@@
However, DMDSpeech raises important ethical concerns. Our model has demonstrated the ability to generate speech with higher perceived similarity to the prompt than real utterances by the same speaker, as judged by both human listeners and speaker verification systems. This highlights limitations in current speaker verification models and presents risks of misuse, such as deepfake generation. To mitigate these risks, more advanced speaker verification techniques capable of distinguishing synthetic from real speech are necessary, alongside robust watermarking to identify synthesized audio. Clear ethical guidelines and legal frameworks will also be crucial to prevent abuse in sensitive areas.
We also observed that while DMDSpeech benefits from fast sampling and direct metric optimization, this comes with a trade-off in speech diversity. The reduction in diversity, which arises from prioritizing sampling speed and quality, warrants further investigation and improvement. Scaling the model with larger datasets and incorporating diverse languages may help mitigate this trade-off.
@@ C.1 DAC Variational Autoencoder
C.2 DMDSpeech
In this section, we present the implementation details of our DMDSpeech model, including the noise schedule, gradient calculation of DMD loss, detailed architecture, and sampling algorithm.
@@ C.2.3 Detailed Architecture
Table 8: DMDSpeech DiT model parameters.
@@ C.2.4 DMD Sampling
Our sampling algorithm of the student (DMDSpeech) is similar to that of the consistency model (Song et al., 2023). The sampling procedure is outlined in Algorithm 1.

### v2

@@ 4.4 Ablation Study
Effects of Distribution Matching Distillation Using a single sampling step resulted in significantly degraded performance compared to the full DMOSpeech model. While increasing to four sampling steps improved naturalness and sound quality to approach the teacher model’s level, speaker similarity remained significantly lower. Interestingly, the speaker verification model’s SIM score showed only a slight decrease, suggesting a phenomenon we term mode shrinkage (Figure 2), where distillation emphasizes high-probability regions of the data distribution. This focus can result in a more generic speaker profile, reducing perceived uniqueness in the prompt speaker’s voice, while maintaining global speaker features as reflected in the SIM score. To address this, we introduced speaker verification loss in this work to better capture the distinct characteristics of the prompt speaker.
@@
Table 4: Ablation study comparing our proposed model with different conditions. MOS-N, MOS-Q, SMOS-V, and SMOS-S are reported as mean ($\pm$ standard error). Models with statistically significant differences ($p<0.05$) compared to DMOSpeech are marked with one asterisk (*). Additional evaluation results on emotion reflection are presented in Table 7.
Model MOS-N MOS-Q SMOS-V SMOS-S WER SIM $\text{CV}_{f_{0}}$ Teacher (N=128) 4.22 ($\pm$ 0.04) 4.40 ($\pm$ 0.03) 4.03 ($\pm$ 0.04) 3.87 ($\pm$ 0.04) 9.51 0.55 0.70 DMD 2 only (N=1) 3.11 ($\pm$ 0.05) 2.99 ($\pm$ 0.05) 2.57 ($\pm$ 0.05) 2.74 ($\pm$ 0.05) 5.93 0.42 0.68 DMD 2 only (N=4) 4.19 ($\pm$ 0.03) 4.43 ($\pm$ 0.04) 3.69 ($\pm$ 0.05) 3.62 ($\pm$ 0.05) 5.67 0.53 0.61 $+\mathcal{L}_{\text{CTC}}$ only 4.25 ($\pm$ 0.04) 4.42 ($\pm$ 0.03) 3.73 ($\pm$ 0.05) 3.62 ($\pm$ 0.05) 1.79 0.55 0.57 $+\mathcal{L}_{\text{SV}}$ only 4.07 ($\pm$ 0.04) 4.33 ($\pm$ 0.03) 4.35 ($\pm$ 0.04) 4.15 ($\pm$ 0.04) 6.62 0.70 0.61 DMOSpeech (N=4) 4.27 ($\pm$ 0.03) 4.45 ($\pm$ 0.03) 4.35 ($\pm$ 0.03) 4.16 ($\pm$ 0.03) 1.94 0.69 0.58 B. S. $96\rightarrow 16$ 4.20 ($\pm$ 0.04) 4.30 ($\pm$ 0.03) 4.27 ($\pm$ 0.04) 4.11 ($\pm$ 0.04) 3.38 0.67 0.60
@@ 5 Conclusions
In this work, we presented DMOSpeech, a distilled diffusion-based text-to-speech (TTS) model based on prompt continuation. By employing distribution matching distillation (DMD), our model generates high-quality speech in just 4 steps and enables direct metric optimization. Through speaker verification (SV) and connectionist temporal classification (CTC) losses, DMOSpeech significantly improves speaker similarity and text-speech alignment, outperforming several state-of-the-art baselines.
@@
However, DMOSpeech raises important ethical concerns. Our model has demonstrated the ability to generate speech with higher perceived similarity to the prompt than real utterances by the same speaker, as judged by both human listeners and speaker verification systems. This highlights limitations in current speaker verification models and presents risks of misuse, such as deepfake generation. To mitigate these risks, more advanced speaker verification techniques capable of distinguishing synthetic from real speech are necessary, alongside robust watermarking to identify synthesized audio. Clear ethical guidelines and legal frameworks will also be crucial to prevent abuse in sensitive areas.
We also observed that while DMOSpeech benefits from fast sampling and direct metric optimization, this comes with a trade-off in speech diversity. The reduction in diversity, which arises from prioritizing sampling speed and quality, warrants further investigation and improvement. Scaling the model with larger datasets and incorporating diverse languages may help mitigate this trade-off.
@@ C.1 DAC Variational Autoencoder
C.2 DMOSpeech
In this section, we present the implementation details of our DMOSpeech model, including the noise schedule, gradient calculation of DMD loss, detailed architecture, and sampling algorithm.
@@ C.2.3 Detailed Architecture
Table 8: DMOSpeech DiT model parameters.
@@ C.2.4 DMD Sampling
Our sampling algorithm of the student (DMOSpeech) is similar to that of the consistency model (Song et al., 2023). The sampling procedure is outlined in Algorithm 1.